OpenStudy (anonymous):

The perimeter of a rectangular painting is 328 centimeters. If the width of the painting is 68 centimeters, what is its length?

OpenStudy (noseboy908):

Do you know the formula for perimeter?

OpenStudy (anonymous):

No :/ need helpp

OpenStudy (noseboy908):

ok, it's P= 2L+2W

OpenStudy (anonymous):

The formula is: 2L + 2W = P (length times 2) + (width x 2) = Permineter So plug in your values. you know P is 328 and w is 68

OpenStudy (anonymous):

2L + 2(68) = 328 and solve for L

OpenStudy (anonymous):

The rest is basic algebra.

OpenStudy (anonymous):

From 2L +2(68) =328, 2L = 328-136 2L = 192 L = 96 right??

OpenStudy (anonymous):

2*68=136. 328-136=192. 192/2=96. So yes, the length of the painting is 96 centimeters.